# Gain characterization of LGAD sensors with beta particles and 28-MeV protons

### v2

@@ 3.1 Beta source
The tandem Van de Graaff Facility at BNL consists of two 15-MV electrostatic accelerators capable of delivering continuous or high-intensity pulsed ion beams to experimental chambers in a wide range of ion species at various energies [18]. For this experiment a beam of 28-MeV protons was used with a flux in the range 1-5 protons/(s cm2). Data were recorded for between 5 minutes and 40 minutes per working point. At least 10,000 events were collected for diodes and for the LGAD at bias voltages below 200 V, and at least 80,000 events were collected for the LGAD at higher voltages. The FNAL readout board with the sensors was placed inside a vacuum chamber that is located at the end of the beam line. The board was connected through a thermal glue to a cold chuck that kept the temperatures stable at about C using a chiller that ran with a fluid consisting of a mix of water and glycol. High voltage to the devices under test was provided by a 16-channel CAEN multi-channel power supply unit which was placed outside the vacuum chamber and controlled remotely by the operators in a control-room through a custom-made interface based on Python drivers.
@@ 6 Simulation
Figure 12 shows the predicted pulses for a diode and an LGAD as generated by the two different particle beams when the sensors are biased at 215 V. In Table 1, the amplitude, the pulse area, the FWHM and the rise times for the LGAD and the diode are extracted from the generated pulses with the bias voltage set at 215 V. The effects of the electronics on pulses are simulated using a first-order low-pass filter response function for the two amplification stages of the readout board and the one of the oscilloscope, using the readout board and oscilloscope parameters presented in Section 2.222For simulating the two amplification stages of the readout board and the one of the oscilloscope, a transfer function of the circuit of the form was used in both cases, where represents the Laplace variable (with being the radian frequency of the sinusoidal signal) and represents the circuit time constant. For the amplification stages, the time constant was set to , with and pF, with the latter accounting for the detector and amplifier capacitance. For the oscilloscope stage, the time constant was set to , with GHz. To simulate the total signal shaping effect, the function for the two amplification stages and the oscilloscope are multiplied together with their respective values, while for the signal normalization, the value of the amplifier gain squared was used, i.e. (see Appendix B). The simulation of the electronics shows qualitatively the change of the sensor performance parameters and is not meant to provide a quantitative assessment of such effect. Figures 13 and 14 show the values of the pulse amplitude, area, FWHM and rise time as functions of the bias voltage, simulated pre- and post-electronics for an LGAD and a diode, respectively, when impinged by a MIP and a 28-MeV proton. The effect of the electronics has a significant impact on the values of FWHM and rise time. The signals generated by the 28-MeV beam have amplitudes 10 times (diode) and about 3 times (LGAD) higher than those generated by a MIP. The values of FWHM and rise time of the LGAD and the diode are very similar in the beams of 28-MeV protons and beta-particles, when the effect of the electronics is included in the simulation. Prior to including the effect of the electronics in the LGAD signal, while the FWHM values decrease for increasing values of bias voltage, the FWHM ratios of 28-MeV protons to beta-particles increase to at 215 V. Conversely, prior to including the effect of the electronics, the rise-time is greater in a beta-particle beam than in a 28-MeV proton beam by a factor close to 2 at 215 V. Table 1 also shows the gains of an LGAD, calculated with the two methods: as the ratio of amplitudes and of the pulse areas in the LGAD and in the diode. With both methods the simulation predicts a gain suppression close to a factor 3, in a 28-MeV proton beam compared to a MIP beam. It is also noteworthy that the two methods provide estimates of gains that differ by a factor of about 3, when the effects of the electronics are not simulated, with the area method providing the largest gain values. The difference becomes far smaller when the effects of the electronics are included.
@@
Amplitude [A] / [mV] | 9.41 (2.12) / 0.309 (0.028) | 31.6 (21.2) / 1.17 (0.281) |
Area [A ns] / [mV ns] | 4.80 (0.420) / 0.476 (0.0415) | 18.4 (4.20) / 1.82 (0.416) |
FWHM [ns] | 0.532 (0.190) / 1.38 (1.33) | 0.599 (0.190) / 1.39 (1.33) |
Rise time [ps] | 232 (12.2) / 341 (293) | 124 (12.2) / 353 (294) |
Gain by Amplitude Method | 4.44 / 11.0 | 1.49 / 4.20 |
Gain by Area Method | 11.4 / 11.4 | 4.37 / 4.40 |
@@ 7 Comparison of results
Figure 15 shows the experimental results for the ratios of amplitude, area, FWHM and rise time. The ratios of pulse amplitudes and areas are very similar. When the sensor is under a high bias voltage, i.e. at or greater than 100 V, the amplitudes and areas in a 28-MeV proton beam are in the range of 5-8 times larger than in a beta-particle beam, in agreement with the simulation that predicts the ratios to be approximately 3-8, regardless of whether the effects of the electronics are included or not. For bias voltages V, the values of FWHM and rise time are stable as functions of bias voltage and very close to 1, in agreement with the simulation, with and without the effects of the electronics. Similarly, the speed of the signal (rise time) is consistent in the two beam types in data, in agreement with the simulation when the effects of the electronics are included, whereas, without the effect of the electronics, the simulation predicts it to be faster in the 28-MeV proton beam by a factor of approximately 2 at 215 V and approximately 10 or greater for lower values of bias voltage.

### v3

@@ 3.1 Beta source
The tandem Van de Graaff Facility at BNL consists of two 15-MV electrostatic accelerators capable of delivering continuous or high-intensity pulsed ion beams to experimental chambers in a wide range of ion species at various energies [18]. For this experiment a beam of 28-MeV protons was used with a flux in the range 1-5 protons/(s cm2). Data were recorded for between 5 minutes and 40 minutes per working point. At least 10,000 events were collected for diodes and for the LGAD at bias voltages below 200 V, and at least 80,000 events were collected for the LGAD at higher voltages. The FNAL readout board with the sensors was placed inside a vacuum chamber that is located at the end of the beam line. The board was connected through a thermal glue to a cold chuck that kept the temperatures stable at about C using a chiller that ran with a fluid consisting of a mix of water and glycol. High voltage to the devices under test was provided by a Keithley 2410 as a power supply unit which was placed outside the vacuum chamber and controlled remotely by the operators in a control-room through a custom-made interface based on Python drivers.
@@ 6 Simulation
Figure 12 shows the predicted pulses for a diode and an LGAD as generated by the two different particle beams when the sensors are biased at 215 V. In Table 1, the amplitude, the pulse area, the FWHM and the rise times for the LGAD and the diode are extracted from the generated pulses with the bias voltage set at 215 V. The effects of the electronics on pulses are simulated using a first-order low-pass filter response function for the two amplification stages of the readout board and the one of the oscilloscope, using the readout board and oscilloscope parameters presented in Section 2.222For simulating the two amplification stages of the readout board and the one of the oscilloscope, a transfer function of the circuit of the form was used in both cases, where represents the Laplace variable (with being the radian frequency of the sinusoidal signal) and represents the circuit time constant. For the amplification stages, the time constant was set to , with and pF, with the latter accounting for the detector and amplifier capacitance. For the oscilloscope stage, the time constant was set to , with GHz. To simulate the total signal shaping effect, the function for the two amplification stages and the oscilloscope are multiplied together with their respective values, while for the signal normalization, the value of the amplifier gain squared was used, i.e. (see Appendix B). The simulation of the electronics shows qualitatively the change of the sensor performance parameters and is not meant to provide a quantitative assessment of such effect. Figures 13 and 14 show the values of the pulse amplitude, area, FWHM and rise time as functions of the bias voltage, simulated pre- and post-electronics for an LGAD and a diode, respectively, when impinged by a MIP and a 28-MeV proton. The effect of the electronics has a significant impact on the values of FWHM and rise time. The signals generated by the 28-MeV beam have amplitudes 10 times (diode) and about 3.7 times (LGAD) higher than those generated by a MIP. The values of FWHM and rise time of the LGAD and the diode are very similar in the beams of 28-MeV protons and beta-particles, when the effect of the electronics is included in the simulation. Prior to including the effect of the electronics in the LGAD signal, while the FWHM values decrease for increasing values of bias voltage, the FWHM ratios of 28-MeV protons to beta-particles increase to at 215 V. Conversely, prior to including the effect of the electronics, the rise-time is greater in a beta-particle beam than in a 28-MeV proton beam by a factor close to 2 at 215 V. Table 1 also shows the gains of an LGAD, calculated with the two methods: as the ratio of amplitudes and of the pulse areas in the LGAD and in the diode. With both methods the simulation predicts a gain suppression close to a factor 3, in a 28-MeV proton beam compared to a MIP beam. It is also noteworthy that the two methods provide estimates of gains that differ by a factor of about 3, when the effects of the electronics are not simulated, with the area method providing the largest gain values. The difference becomes far smaller when the effects of the electronics are included.
@@
| Amplitude [A] / [mV] | 9.4 (2.1) / 23 (2.3) | 32 (21) / 84 (22) |
| Area [A ns] / [mV ns] | 4.8 (0.42) / 24 (2.1) | 18 (4.2) / 91 (21) |
| FWHM [ns] | 0.53 (0.19) / 0.91 (0.79) | 0.60 (0.19) / 0.94 (0.80) |
| Rise time [ps] | 230 (12) / 250 (170) | 120 (12) / 270 (170) |
| Gain by Amplitude Method | 4.4 / 9.9 | 1.5 / 3.8 |
| Gain by Area Method | 11 / 11 | 4.3 / 4.3 |
@@ 7 Comparison of results
Figure 15 shows the experimental results for the ratios of amplitude, area, FWHM and rise time. The ratios of pulse amplitudes and areas are very similar. When the sensor is under a high bias voltage, i.e. at or greater than 100 V, the amplitudes and areas in a 28-MeV proton beam are in the range of 5-8 times larger than in a beta-particle beam, in agreement with the simulation that predicts the ratios to be approximately 3-10, regardless of whether the effects of the electronics are included or not. For bias voltages V, the values of FWHM and rise time are stable as functions of bias voltage and very close to 1, in agreement with the simulation, with and without the effects of the electronics. Similarly, the speed of the signal (rise time) is consistent in the two beam types in data, in agreement with the simulation when the effects of the electronics are included, whereas, without the effect of the electronics, the simulation predicts it to be faster in the 28-MeV proton beam by a factor of approximately 2 at 215 V and approximately 10 or greater for lower values of bias voltage.